What is the power output of a cyclist moving at a steady speed of 5ms-1 along a level road against resistance of 20N
1
Expert's answer
2020-11-06T10:13:34-0500

"\text{according to Newton's 3 law, the action force is equal in magnitude and opposite}" "\text{in direction to the reaction force}"

"F=20;\ P= F*V*\cos(\alpha)\ ;\alpha= 0"

"P= 20*5*1 =100 \text{ watt}"

Answer: "P= 100 \text{ watt}"